# The Nondecreasing Rank

### v1

@@ Abstract
Keywords: contingency table; hierarchical alternating least squares, Möbius inversion, nonnegative matrix factorization; partially ordered set; polytope; simplicial cone; tensor; rank.
@@ 1 Introduction
Nonnegative matrix factorizations have a statistical motivation as they also arise in the study of contingency tables as mixture models of independent random variables [21, Ch 4]. One strand of literature within the field of algebraic statistics has been to understand the semialgebraic geometry of such mixture models. Despite being simple to formulate, the geometry of NMF is involved, even in the rank-$3$ case [22, 32].
@@
An overview of this paper is as follows: Section 3 provides background material and Section 4 provides further introduction and motivation for the concept of the ND rank. Sections 5 and 6 address the questions of when exact ND factorizations exist and when they can be reframed as nonnegative factorizations up to a linear change in coordinates. The geometry of order cones is also explicated here. Section 7 provides exact expressions for the maximum ND rank in certain cases. It is shown that knowing the maximum ND rank is enough to fully determine the typical ND ranks. Section 8 shows that if the rank of a matrix is one or two and it has a finite ND rank, then the ND rank is also equal to one or two respectively. Section 9 shows that the border ND rank is equivalent to the ND rank, a fact which is relevant for optimization. Section 10 proposes an ND hierarchical least squares optimization algorithm. The paper concludes with an application section illustrating how ND factorizations are fruitful for uncovering structure in data.
@@ 2 Notation
Some of the notation used in this work is listed here for easy reference. Partially ordered sets (posets) are represented by $\mathcal{P}$, the order cone associated with this poset is denoted by $\mathcal{C}(\mathcal{P})$, and the set of tensors with ND rank at most $r$ is $\mathcal{N}_{\leq r}$. The dual of a cone $\mathcal{C}$ is $\mathcal{C}^{*}$. A curly inequality $\prec$ represents inequality with respect to a poset while $x\lessdot y$, means that the element $y$ covers $x$ in the poset. $\mathbb{R}^{p}_{+}$ is equal to all vectors in $\mathbb{R}^{p}$ with nonnegative entries and ${\mathbf{e}}_{i}$ is the $i$th standard basis vector. Vectors, matrices, and tensors are all displayed in bold font, while scalars are not. Given any finite dimensional vector spaces $V_{1},\ldots,V_{k}$ and sets $S_{i}\subseteq V_{i}$, $\otimes_{i=1}^{k}V_{i}$ is the tensor product space of the $V_{i}$ and $\otimes_{i=1}^{k}S_{i}$ is the subset $\cup_{r=1}^{\infty}\{\sum_{i=1}^{r}\otimes_{j=1}^{k}{\mathbf{s}}^{(ij)}:{\mathbf{s}}^{(ij)}\in S_{j}\}$ of this tensor product space. We will use $k$ to refer to the number of factors in the tensor product and take $p_{i}\coloneq\text{dim}(V_{i})$, where usually $V_{i}=\mathbb{R}^{\mathcal{P}_{j}}$ with $\mathcal{P}_{j}$ a poset containing $p_{j}$ elements. Shorthand for the set $\{1,\ldots,p\}$ is $[p]$. The standard inner product between two order-$k$ tensors is $\langle{\mathbf{S}},{\mathbf{T}}\rangle=\sum_{i_{1}=1}^{p_{1}}\cdots\sum_{i_{k}=1}^{p_{k}}S_{i_{1}\ldots i_{k}}T_{i_{1}\ldots i_{k}}$. This induces the Frobenius norm $\|{\mathbf{T}}\|_{F}^{2}=\langle{\mathbf{T}},{\mathbf{T}}\rangle$.
@@ 3 Preliminaries on Low-rank Tensors and Convex Geometry
where every $({\mathbf{v}}^{(j)})^{\intercal}=(v_{1}^{(j)},\ldots,v_{p_{j}}^{(j)})$ is a vector in $\mathbb{R}^{p_{j}}$. The tensor product notation $\otimes$ is used to denote the above rank-one structure: ${\mathbf{T}}=\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}$. The real (CP [31, Sec 3]) rank of a tensor is defined as
$\text{rank}({\mathbf{T}})=\min\bigg{\{}r:{\mathbf{T}}=\sum_{i=1}^{r}\otimes_{j=1}^{k}{\mathbf{v}}^{(ij)}\bigg{\}}.$ (1)
That is, it is the smallest number of rank-one tensors required to express $\mathbf{T}$} as a sum of rank-one tensors. Whenever rank is referred to in this work it is taken to mean rank over the real numbers. The nonnegative rank [28] places extra nonnegativity constraints on the vectors ${\mathbf{v}}^{(ij)}$ and is defined as
@@
In general, the nonnegative rank of a tensor may exceed its rank [16].
@@
A (convex) cone $\mathcal{C}\subset\mathbb{R}^{p}$ satisfies the property that for all ${\mathbf{x}},{\mathbf{y}}\in\mathcal{C}$ and $\alpha,\beta\geq 0$ the vector $\alpha{\mathbf{x}}+\beta{\mathbf{y}}$ is in $\mathcal{C}$. The conical hull of the vectors ${\mathbf{v}}_{1},\ldots,{\mathbf{v}}_{m}$ is defined as $\{\sum_{i=1}^{m}\alpha_{i}{\mathbf{v}}_{i}:\alpha_{i}\geq 0\}$. Any cone that can be expressed in this way is called polyhedral with the representation referred to as a vertex or $\mathcal{V}$-representation. Polyhedral cones can also equivalently be expressed in a halfspace or $\mathcal{H}$-representation, where a cone is defined by finitely many halfspace constraints: $\{{\mathbf{x}}:{\mathbf{a}}_{i}^{\intercal}{\mathbf{x}}\geq 0,i=1,\ldots,l\}$. The dual cone $\mathcal{C}^{*}$ is closely related to the $\mathcal{H}$-representation as it is defined as the closed, convex set $\mathcal{C}^{*}=\{{\mathbf{a}}:{\mathbf{a}}^{\intercal}{\mathbf{x}}\geq 0,\forall{\mathbf{x}}\in\mathcal{C}\}$. A cone is said to be pointed if $\mathcal{C}\cap(-\mathcal{C})=\emptyset$, solid if it has a non-empty interior, and proper if it is pointed, solid, and closed. Polyhedral cones are pointed if and only if the ${\mathbf{a}}_{i}$ appearing in the $\mathcal{H}$-representation span $\mathbb{R}^{p}$, are solid if and only if the ${\mathbf{v}}_{i}$ appearing in the $\mathcal{V}$-representation span $\mathbb{R}^{p}$, and are always closed. Whenever $\mathcal{C}$ is polyhedral or pointed the dual cone $\mathcal{C}^{*}$ inherits these properties respectively.
@@ 4 The Nondecreasing Rank
If a concise representation of this data was sought, a simple assumption would be that the dose-response functions $f_{i}:\{0,100,200,400\}\rightarrow[0,100]$ for each type of selenium satisfy $f^{(i)}=c_{i}f$ for some nondecreasing function $f$ and scalars $c_{3}\geq c_{1},c_{2}\geq 0$, with the latter assumption coming from the observation that the third type of selenium appears to be most harmful. Formally, this representation amounts to assuming that the dose-response matrix ${\mathbf{T}}\approx{\mathbf{c}}{\mathbf{f}}^{\intercal}$ is rank-one and that the nonnegative vectors $\mathbf{c}$} and $\mathbf{v}$} are congruous with the respective orderings Type I, $I\text{Type I, Type II}\leq\text${Type III}, and $0\leq 100\leq 200\leq 400$ that are imposed on the row and column variables of the matrix. If $\mathbf{c}$} is scaled so that $\tfrac{1}{3}(c_{1}+c_{2}+c_{3})=1$ then the vector $\mathbf{f}$} could be interpreted as the average dose-response curve across all three types of selenium, while $\mathbf{c}$} reflects the multiplicative noxious effects of the different varieties of selenium.
@@
The nondecreasing rank generalizes the idea from the dose-response example to tensors defined on index sets that are products of partially ordered sets (posets). For $i=1,\ldots,k$ let $\mathcal{P}_{j}$ be a partially ordered set with $p_{j}$ elements. The product $\mathcal{P}=\times_{j=1}^{k}\mathcal{P}_{j}$ is a partially ordered set consisting of $k$-tuples of elements where $(x_{1},\ldots,x_{k})\preceq(y_{1},\ldots,y_{k})$ if and only if $x_{j}\preceq y_{j}$ in $\mathcal{P}_{j}$ for every $j$. In the selenium example above, $\mathcal{P}_{1}=\{\mathrm{Type\,I,Type\,II,Type\,III}\}$ is a poset and $\mathcal{P}_{2}=\{0,100,200,400\}$ is also a poset under the natural ordering. In $\mathcal{P}_{1}\times\mathcal{P}_{2}$ we have that $(\mathrm{Type\,II},300)\preceq(\mathrm{Type\,III},400)$ but $(\mathrm{Type\,III},0)\npreceq(\mathrm{Type\,I},400)$. Figures 1 and 2 illustrate Hasse diagrams of the partial orderings in $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, and $\mathcal{P}_{1}\times\mathcal{P}_{2}$. In a Hasse diagram an arrow is drawn from an element $x$ to an element $y$ if and only if $y$ covers $x$, meaning that $x\prec y$ and there does not exist an element $z$ with $x\prec z\prec y$. This is denoted by $x\lessdot y$. For instance, there is an arrow from 0 to $100$ but not from 0 to $200$ since $100$ lies in between 0 and $200$. The poset $\mathcal{P}_{2}$ has the additional property of being totally ordered, also called a chain: for any two elements $x,y\in\mathcal{P}_{2}$ either $x\preceq y$ or $y\preceq x$. The poset $\mathcal{P}_{1}$ is not a chain since Type I is neither less than or greater than Type II.
@@
Figure 1: The respective Hasse diagrams of $\{\text{Type I,\, Type II,\, Type III}\}$ and $\{0,100,200,400\}$.
@@
From the perspective that finite-dimensional tensors are functions defined on a Cartesian product of indices, we can consider tensors as functions defined on $\mathcal{P}$. We are primarily interested in tensors that also respect the product ordering of $\mathcal{P}$.
Definition 1 (Order cone and order polytope of functions).
A real-valued function ${\mathbf{f}}\in\mathbb{R}^{\mathcal{P}}$ with an index set $\mathcal{P}$ that is any finite poset is nondecreasing if $f_{x}\leq f_{y}$ whenever $x\preceq y$ in $\mathcal{P}$. The order cone $\mathcal{C}(\mathcal{P})$ consists of all nonnegative, nondecreasing functions on $\mathcal{P}$, while the order polytope $\mathcal{O}(\mathcal{P})$ consists of all nonnegative tensors with $f_{x}\leq 1$ for all $x\in\mathcal{P}$ [50].
The order cone is a proper, polyhedral cone that has the $\mathcal{H}$-representation
Pairing a monotonicity requirement with a low-rank factorization we now introduce the central notion of interest in this work:
Definition 2 (Nondecreasing (ND) rank).
Let $\mathcal{P}=\times_{j=1}^{k}\mathcal{P}_{j}$ be a product poset and consider the vector space $\mathbb{R}^{\mathcal{P}_{1}\times\cdots\times\mathcal{P}_{k}}$ of functions defined on $\mathcal{P}$. The set of tensors with nondecreasing rank of at most $r$ is defined as
$\mathcal{N}_{\leq r}(\mathcal{P})=\bigg{\{}{\mathbf{T}}\in\mathcal{C}(\mathcal{P}):{\mathbf{T}}=\sum_{i=1}^{r}\otimes_{j=1}^{k}{\mathbf{v}}^{(ij)},\;\;{\mathbf{v}}^{(ij)}\in\mathcal{C}(\mathcal{P}_{j}),\;\forall i,j\bigg{\}}.$ (3)
Any such decomposition ${\mathbf{T}}=\sum_{i=1}^{r}\otimes_{j=1}^{k}{\mathbf{v}}^{(ij)}$ as in (3) is referred to as a nondecreasing decomposition or factorization. The set of tensors with a nondecreasing rank of $r$ is defined as $\mathcal{N}_{r}(\mathcal{P})=\mathcal{N}_{\leq r}(\mathcal{P})\backslash\mathcal{N}_{\leq r-1}(\mathcal{P})$. The nondecreasing rank of $\mathbf{T}$} is
$\mathrm{NDrank}({\mathbf{T}})=\min\{r:{\mathbf{T}}\in\mathcal{N}_{\leq r}(\mathcal{P})\},$ (4)
which is infinite if $\mathbf{T}$} does not possess an ND rank-$r$ decomposition for any $r$. The set of finite ND rank tensors is denoted by $\mathcal{N}_{<\infty}(\mathcal{P})$. The poset $\mathcal{P}$ will generally be dropped from the notation $\mathcal{N}_{\leq r}(\mathcal{P}),\mathcal{N}_{<\infty}(\mathcal{P})$.
The ND rank refines the nonnegative rank (2) by adding extra constraints, implying that $\mathrm{rank}_{+}({\mathbf{T}})\leq\mathrm{NDrank({\mathbf{T}}})$. As tensor products are multiplicative, the nonnegativity constraint in the definition of the order cone has an important role in ensuring that monotonicity is preserved under tensor products. Specifically, if the vectors ${\mathbf{v}}^{(ij)}$ were only required to be nondecreasing, but possibly had negative entries, the tensor $\otimes_{i=1}^{k}{\mathbf{v}}^{(ij)}$ would not necessarily be nondecreasing over $\mathcal{P}$.
We note that the terminology of the monotone rank has appeared in the literature [2], although it is synonymous with the nonnegative rank over an ordered field. Other very general notions of rank that encompass the ND rank as a special case are that of the X-rank [42] and the atomic cone rank [23].

### v2

@@ Abstract
Keywords: contingency table; hierarchical alternating least squares; Möbius inversion; nonnegative matrix factorization; partially ordered set; polytope; simplicial cone; tensor; rank.
@@ 1 Introduction
Nonnegative matrix factorizations have a statistical motivation as they also arise in the study of contingency tables as mixture models of independent random variables [22, Ch 4]. One strand of literature within the field of algebraic statistics has been to understand the semialgebraic geometry of such mixture models. Despite being simple to formulate, the geometry of NMF is involved, even in the rank-$3$ case [23, 33].
@@
An overview of this paper is as follows: Section 3 provides background material and Section 4 provides further introduction and motivation for the concept of the ND rank. Sections 5 and 6 address the questions of when exact ND factorizations exist and when they can be reframed as nonnegative factorizations up to a linear change in coordinates. The geometry of order cones is also explicated here. Section 7 provides exact expressions and bounds for the maximum ND rank in certain cases. It is shown that knowing the maximum ND rank is enough to fully determine the typical ND ranks. Section 8 shows that if the rank of a matrix is one or two and it has a finite ND rank, then the ND rank is also equal to one or two respectively. Section 9 shows that the border ND rank is equivalent to the ND rank, a fact which is relevant for optimization. Section 10 proposes an ND hierarchical least squares optimization algorithm. The paper concludes with an application section illustrating how ND factorizations are fruitful for uncovering structure in data.
@@ 2 Notation
Some of the notation used in this work is listed here for easy reference. Partially ordered sets (posets) are represented by $\mathcal{P}$, the order cone associated with this poset is denoted by $\mathcal{C}(\mathcal{P})$, and the set of tensors with ND rank at most $r$ is $\mathcal{N}_{\leq r}$. The dual of a cone $\mathcal{C}$ is $\mathcal{C}^{*}$. A curly inequality $\prec$ represents inequality with respect to a poset while $x\lessdot y$, means that the element $y$ covers $x$ in the poset. $\mathbb{R}^{p}_{+}$ is equal to all vectors in $\mathbb{R}^{p}$ with nonnegative entries and ${\mathbf{e}}_{i}$ is the $i$th standard basis vector. Vectors, matrices, and tensors are all displayed in bold font, while scalars are not. Given any finite-dimensional vector spaces $V_{1},\ldots,V_{k}$ and sets $S_{i}\subseteq V_{i}$, $\otimes_{i=1}^{k}V_{i}$ is the tensor product space of the $V_{i}$ and $\otimes_{i=1}^{k}S_{i}$ is the subset $\cup_{r=1}^{\infty}\{\sum_{i=1}^{r}\otimes_{j=1}^{k}{\mathbf{s}}^{(ij)}:{\mathbf{s}}^{(ij)}\in S_{j}\}$ of this tensor product space. We will use $k$ to refer to the number of factors in the tensor product and take $p_{i}\coloneq\text{dim}(V_{i})$, where usually $V_{i}=\mathbb{R}^{\mathcal{P}_{j}}$ with $\mathcal{P}_{j}$ a poset containing $p_{j}$ elements. Shorthand for the set $\{1,\ldots,p\}$ is $[p]$. The standard inner product between two order-$k$ tensors is $\langle{\mathbf{S}},{\mathbf{T}}\rangle=\sum_{i_{1}=1}^{p_{1}}\cdots\sum_{i_{k}=1}^{p_{k}}S_{i_{1}\ldots i_{k}}T_{i_{1}\ldots i_{k}}$. This induces the Frobenius norm $\|{\mathbf{T}}\|_{F}^{2}=\langle{\mathbf{T}},{\mathbf{T}}\rangle$.
@@ 3 Preliminaries on Low-rank Tensors and Convex Geometry
where every $({\mathbf{v}}^{(j)})^{\intercal}=(v_{1}^{(j)},\ldots,v_{p_{j}}^{(j)})$ is a vector in $\mathbb{R}^{p_{j}}$. The tensor product notation $\otimes$ is used to denote the above rank-one structure: ${\mathbf{T}}=\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}$. The real (CP [32, Sec 3]) rank of a tensor is defined as
$\text{rank}({\mathbf{T}})=\min\bigg\{r:{\mathbf{T}}=\sum_{i=1}^{r}\otimes_{j=1}^{k}{\mathbf{v}}^{(ij)}\bigg\},$ (1)
where every ${\mathbf{v}}^{(ij)}\in\mathbb{R}^{p_{j}}$. That is, it is the smallest number of rank-one tensors required to express $\mathbf{T}$} as a sum of rank-one tensors. Whenever rank is referred to in this work it is taken to mean rank over the real numbers. The nonnegative rank [29] places extra nonnegativity constraints on the vectors ${\mathbf{v}}^{(ij)}$ and is defined as
@@
In general, the nonnegative rank of a tensor may exceed its rank [17, 55].
@@
A (convex) cone $\mathcal{C}\subset\mathbb{R}^{p}$ satisfies the property that for all ${\mathbf{x}},{\mathbf{y}}\in\mathcal{C}$ and $\alpha,\beta\geq 0$ the vector $\alpha{\mathbf{x}}+\beta{\mathbf{y}}$ is in $\mathcal{C}$. The conical hull of the vectors ${\mathbf{v}}_{1},\ldots,{\mathbf{v}}_{m}$ is defined as $\{\sum_{i=1}^{m}\alpha_{i}{\mathbf{v}}_{i}:\alpha_{i}\geq 0\}$. Any cone that can be expressed in this way is called polyhedral with the representation referred to as a vertex or $\mathcal{V}$-representation. Polyhedral cones can also equivalently be expressed in a halfspace or $\mathcal{H}$-representation, where a cone is defined by finitely many halfspace constraints: $\{{\mathbf{x}}:{\mathbf{a}}_{i}^{\intercal}{\mathbf{x}}\geq 0,i=1,\ldots,l\}$. The dual cone $\mathcal{C}^{*}$ is closely related to the $\mathcal{H}$-representation as it is defined as the closed, convex set $\mathcal{C}^{*}=\{{\mathbf{a}}:{\mathbf{a}}^{\intercal}{\mathbf{x}}\geq 0,\forall{\mathbf{x}}\in\mathcal{C}\}$. A cone is said to be pointed if $\mathcal{C}\cap(-\mathcal{C})=\{{\mathbf{0}}\}$, solid if it has a non-empty interior, and proper if it is pointed, solid, and closed. Polyhedral cones are pointed if and only if the ${\mathbf{a}}_{i}$ appearing in the $\mathcal{H}$-representation span $\mathbb{R}^{p}$, are solid if and only if the ${\mathbf{v}}_{i}$ appearing in the $\mathcal{V}$-representation span $\mathbb{R}^{p}$, and are always closed. Whenever $\mathcal{C}$ is polyhedral or proper the dual cone $\mathcal{C}^{*}$ inherits these properties respectively.
@@ 4 The Nondecreasing Rank
If a concise representation of this data was sought, a simple assumption would be that the dose-response functions $f_{i}:\{0,100,200,400\}\rightarrow[0,100]$ for each type of selenium satisfy $f^{(i)}=c_{i}f$ for some nondecreasing function $f$ and scalars $c_{3}\geq c_{1},c_{2}\geq 0$, with the latter assumption coming from the observation that the third type of selenium appears to be most harmful. This representation amounts to assuming that the dose-response matrix ${\mathbf{T}}\approx{\mathbf{c}}{\mathbf{f}}^{\intercal}$ is rank-one and that the nonnegative vectors $\mathbf{c}$} and $\mathbf{v}$} are congruous with the respective orderings Type I, $I\text{Type I, Type II}\leq\text${Type III}, and $0\leq 100\leq 200\leq 400$ that are imposed on the row and column variables of the matrix. If $\mathbf{c}$} is scaled so that $\tfrac{1}{3}(c_{1}+c_{2}+c_{3})=1$ then the vector $\mathbf{f}$} could be interpreted as the average dose-response curve across all three types of selenium, while $\mathbf{c}$} reflects the multiplicative, noxious effects of the different varieties of selenium.
Formally, in the selenium example we have associated two partially ordered sets (posets) with the rows and columns of the table, namely $\mathcal{P}_{1}=\{\mathrm{Type\,I,Type\,II,Type\,III}\}$ and $\mathcal{P}_{2}=\{0,100,200,400\}$. The vectors $\mathbf{c}$} and $\mathbf{f}$} can be viewed as functions defined on $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ respectively. To say that both $\mathbf{c}$} and $\mathbf{f}$} respect the given partial orderings is equivalent to saying that these vectors lie in the order cones $\mathcal{C}(\mathcal{P}_{1})$ and $\mathcal{C}(\mathcal{P}_{2})$:
Definition 1 (Order cone and order polytope of functions).
Let $\mathcal{P}$ be a finite poset and take ${\mathbf{f}}\in\mathbb{R}^{\mathcal{P}}$ be a real-valued function defined on this poset. The function $\mathbf{f}$} is said to be nondecreasing if $f_{x}\leq f_{y}$ whenever $x\preceq y$ in $\mathcal{P}$. The order cone $\mathcal{C}(\mathcal{P})$ consists of all nonnegative, nondecreasing functions on $\mathcal{P}$, while the order polytope $\mathcal{O}(\mathcal{P})$ consists of all tensors in $\mathcal{C}(\mathcal{P})$ with $f_{x}\leq 1$ for all $x\in\mathcal{P}$ [51].
An order cone is a proper, polyhedral cone that has the $\mathcal{H}$-representation
It may be the case that a rank-one factorization of the form ${\mathbf{c}}{\mathbf{f}}^{\intercal}$ is not sufficient to capture all of the structure inherent to the selenium dose-response matrix; more general low-rank representations are sought. Pairing a monotonicity requirement with a low-rank factorization we now introduce the central notion of interest in this work:
Definition 2 (Nondecreasing (ND) rank).
Let ${\mathbf{T}}\in\mathbb{R}^{\mathcal{P}_{1}\times\cdots\times\mathcal{P}_{k}}$ be a tensor defined on the Cartesian product index set $\mathcal{P}\coloneq\mathcal{P}_{1}\times\cdots\times\mathcal{P}_{k}$, where each $\mathcal{P}_{j}$ is a finite poset. Any decomposition of the form ${\mathbf{T}}=\sum_{i=1}^{r}\otimes_{j=1}^{k}{\mathbf{v}}^{(ij)}$ with ${\mathbf{v}}^{(ij)}\in\mathcal{C}(\mathcal{P}_{j})$ is referred to as a nondecreasing decomposition or factorization. The nondecreasing rank of a tensor $\mathbf{T}$} is the minimum number of rank-one factors needed in a nondecreasing factorization:
$\mathrm{NDrank}({\mathbf{T}})=\min\bigg\{r:{\mathbf{T}}=\sum_{i=1}^{r}\otimes_{j=1}^{k}{\mathbf{v}}^{(ij)},\;\;{\mathbf{v}}^{(ij)}\in\mathcal{C}(\mathcal{P}_{j}),\;\forall i,j\bigg\}.$ (3)
If $\mathbf{T}$} does not possess a nondecreasing factorization then it has an infinite ND rank. The set of all tensors with finite ND rank, ND rank equal to $r$, and ND rank less than or equal to $r$ are respectively denoted by $\mathcal{N}_{<\infty}(\mathcal{P})$, $\mathcal{N}_{r}(\mathcal{P})$, and $\mathcal{N}_{\leq r}(\mathcal{P})$. The poset $\mathcal{P}$ will generally be dropped from this notation.
The ND rank refines the nonnegative rank (2) by possibly adding extra constraints, implying that $\mathrm{rank}_{+}({\mathbf{T}})\leq\mathrm{NDrank({\mathbf{T}}})$. We note that the terminology of monotone rank has appeared in the literature [2], although it is synonymous with the nonnegative rank over an ordered field. Other very general notions of rank that encompass the ND rank as a special case are that of the X-rank [43] and the atomic cone rank [24].
@@
The tensor $\mathbf{T}$} appearing in the definition of ND rank can be viewed as a function on the Cartesian product of posets $\mathcal{P}\coloneq\times_{j=1}^{k}\mathcal{P}_{j}$. The set $\mathcal{P}$ can itself be given the structure of a poset, called the product poset, by defining the ordering with $(x_{1},\ldots,x_{k})\preceq(y_{1},\ldots,y_{k})$ if and only if $x_{j}\preceq y_{j}$ in $\mathcal{P}_{j}$ for every $j$. In the selenium example we have that $(\mathrm{Type\,II},300)\preceq(\mathrm{Type\,III},400)$ but $(\mathrm{Type\,III},0)\npreceq(\mathrm{Type\,I},400)$. A helpful way to visualize posets is by their Hasse diagrams. A Hasse diagram is a directed graph with nodes given by the elements of the poset and arrows drawn from an element $x$ to an element $y$ if and only if $y$ covers $x$, meaning that $x\prec y$ and there does not exist an element $z$ with $x\prec z\prec y$. The covering relation is denoted by $x\lessdot y$. Figures 1 and 2 illustrate the Hasse diagrams of the partial orderings $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, and $\mathcal{P}_{1}\times\mathcal{P}_{2}$ in the selenium example. We see that there is an arrow from 0 to $100$ as $0\lessdot 100$, but not from 0 to $200$ since $100$ lies in between 0 and $200$. The Hasse diagram of the poset $\mathcal{P}_{2}$ is a path and has the additional property of being totally ordered. Totally ordered posets are referred to as chains: for any two elements $x,y\in\mathcal{P}_{2}$ either $x\preceq y$ or $y\preceq x$. The poset $\mathcal{P}_{1}$ is not a chain since Type I is neither less than or greater than Type II.
@@
Figure 1: The respective Hasse diagrams of $\mathcal{P}_{1}=\{\text{Type I,\, Type II,\, Type III}\}$ and $\mathcal{P}_{2}=\{0,100,200,400\}$.
@@
It is worth contemplating why there is a requirement that vectors must be nonnegative in the definition of the order cone, and accordingly in ND factorizations. One motivation for this is that if ${\mathbf{T}}=\sum_{i=1}^{r}\otimes_{j=1}^{k}{\mathbf{v}}^{(ij)}$ is an ND factorization then ${\mathbf{T}}\in\mathcal{C}(\mathcal{P})$ is guaranteed to also respect the order of the product poset $\mathcal{P}$. If we contrarily permitted a nondecreasing ${\mathbf{v}}^{(ij)}$ to contain negative entries the tensor $\otimes_{j=1}^{k}{\mathbf{v}}^{(ij)}$ would not necessarily be nondecreasing over $\mathcal{P}$.